• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name C. C. Chang

    Log entry time 16:05:51 on February23,2002

    Entry number 3758

    keyword=Shift Summary

    08:15 FPP tripped.
    08:37 Called MCC to make hall access to fix gas problem (to increase gas flow
    on FPP left Ch. 4)
    09:15 VDC left bottom tripped.
    09:20 Back to data taking.
    09:40 ROC4 frozen. Rebooted a few times, but failed to restart.
    10:20 Call Ting Chang; also call MCC to make hall access.
    10:55 ROC4 problem fixed (see HALOG # 3704). Call MCC to restore beam.
    11:08 Beam is back and we resume data taking.
    11: 10 FPP left Ch. 4 HV seems stable now after increasing the gas flow rate.
    11:42 FPP left Ch. 4 tripped again!
    11:45 Alarm comes on and off indicating left VDC DISC LVL +5V (power
    supply?) seems to fluctuate; alarm comes on when it is over +5.5V.
    13:20 FPP left Ch. 4 tripped.
    14:00 TS2 hanged, but reset successfully.
    14:47 ROC4 hanged, but resedt successfully this time.
    16:00 Has been a good shift with few minor problems.